eT. For renstors of 75 Ω, 50 Ω. & 100 Ω are connected as shown to a 9 volt source. a) Make a sketch of this circuit (include correct schematic diagrams for ALL components) b) Determine the curreni through the circuit if the switch S is OFF and if the switch S is ON: c) Determine the power dissipated by the circuit if the switch S is OFF and if the switch S is ON:
Answer #1

b) when S is off
R = 100 + 50 = 150
I = V/R = 0.06 A
when S is on
R = 150/2 = 75
I = 0.12 A
c) Power, for S is OFF = VI = 0.54 W
Power, S on = 1.08 W